OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 48, 12, 3, 3/4, ...

OpenStudy (anonymous):

divides by 4..? converges

OpenStudy (turingtest):

to what?

OpenStudy (anonymous):

it either coverges by 64, -4080 or 0

OpenStudy (turingtest):

given the choices, and the fact that the numbers are getting divided each time, which do you think it is?

OpenStudy (anonymous):

would it be 0

OpenStudy (anonymous):

Your sequence is \[\left\{48\left(\frac{1}{4}\right)^k\right\}\] To find the value that the sequence converges to, find \[\lim_{k\to\infty}48\left(\frac{1}{4}\right)^k\]

OpenStudy (anonymous):

Yes, it is 0.

OpenStudy (anonymous):

thanks!